40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Engineering group Lab elog  Not logged in ELOG logo
Message ID: 254     Entry time: Thu Sep 9 14:56:26 2021
Author: Stephen 
Type: Progress 
Category: RTS for COC 
Subject: Anneal of SN1535 at 300C for 10 hours 

SN1535 was annealed between 07 and 08 September 2021 - used the large furnace in Downs 221, atmosphere was air, used glass petri dishes to protect optic (from Gabriele and CRiMe).

Controller parameters:

Ramp up rate = 100 °C per hour
Hold temp = 300 °C
Hold time = 10 hours
Ramp down rate = 100 °C per hour

Witness RTD observed a ~21 °C overshoot and ~14 °C offset from controller RTD. These parameters were consistent with the elog entry ENG_Labs/253. See attached image of the temperature profile (.xlsx file is source document).

Optic is now with Liyuan in RTS for recharacterization after anneal.

Attachment 1: anneal_sn1535_300C_20210910.png  63 kB  Uploaded Mon Sep 13 16:43:28 2021  | Hide | Hide all
anneal_sn1535_300C_20210910.png
Attachment 2: anneal_sn1535_300C_20210910.xlsx  92 kB  Uploaded Mon Sep 13 16:44:45 2021
ELOG V3.1.3-